OpenStudy (anonymous):

315/1575. write the prime fractorization of the numerator and the denominator. then write the fraction in lowest terms.

OpenStudy (watchmath):

\(315=5\cdot 3^2\cdot 7\) and \(1575 =3^2\cdot 5^2\cdot 7\) Then \[315/1575=1/5\]